OpenStudy (anonymous):

Let V be a vector space, with real numbers as scalars, which is equipped with an inner product (, ). (a) Establish the identity ||x + y|| = ||x − y||, where x and y are two orthogonal vectors in V

OpenStudy (anonymous):

It is sufficient to show that the normed squared (ie both sides squared) are equal. ||x +y||^2=(x+y, x+y) the inner product =(x,x) + (x,y) + (y, x) + (y, y) but since x and y orthogonal (x,y)=(y,x)=0 => = (x,x)+(y,y) = ||x||^2 + ||y||^2 Now the RHS is the same process ||x-y||^2=(x-y, x-y) = (x,x) - (x, y) - (y, x) + (y, y) same orthogonality argument =||x||^2+||y||^2 so both sides are equal ||x + y||^2 = ||x-y||^2 implying the identity

OpenStudy (anonymous):

Thanks so much appreciate it!! :)

OpenStudy (anonymous):

A geometric way of thinking about this is that by virtue of the triangle inequality |x+y| <= |x| + |y|, the distance function |x-y| = |y-x| is a metric for V.